OpenStudy (anonymous):

please help! Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. - tan2x + sec2x = 1

OpenStudy (anonymous):

Didn't I just answer this question?

OpenStudy (anonymous):

you didnt explain how to do it

OpenStudy (anonymous):

Well don't delete the question and start a new one, just say so... Use this to rewrite the left side: tan^2x = sin^2x / cos^2x sec^2x = 1 / cos^2x

OpenStudy (anonymous):

so it would be -sin^2x/cos^2x + 1/cos^2x =1 where do I go from here?

OpenStudy (anonymous):

You have a common denominator!

OpenStudy (anonymous):

so -sin^2x +1/cos^2x

OpenStudy (anonymous):

no.. \[\frac{ -\sin^2x }{ \cos^2x } + \frac{ 1 }{ \cos^2x }\] Common denominators means you can add them together

OpenStudy (anonymous):

What do you mean add them together? Isn't that what I just did?

OpenStudy (anonymous):

ohhhh you didn't put brackets you should have written (-sin^2x +1)/cos^2x Very different! Ok good now you can simplify the numerator using the fact that sin^2x + cos^2x + 1

OpenStudy (anonymous):

so am I trying to get rid of the denominator now? which would make it -sin^2x + cos^2x +1?

OpenStudy (anonymous):

How do I finish this?

OpenStudy (anonymous):

cool thanks for answering

OpenStudy (anonymous):

Sorry I had to go. sin^2x + cos^2x =1 cos^2x = 1- sin^2x = -sin^2x +1 substitute this into the fraction (-sin^2x +1)/cos^2x =cos^2x/cos^2x =1